Need some help
Find the equation of a line that contains the point (-2,4)and is parallel to
y=2x-3
thank you for your time.

Answer by fractalier(6550)

Parallel means same slope...so we need a line going thru (-2, 4) with a slope of 2...we use the point-slope formula...
y - y1 = m(x - x1)
y - 4 = 2(x -(-2))
y - 4 = 2(x + 2)
y - 4 = 2x + 4
y = 2x + 8